OpenStudy (anonymous):

g={(0,7),(5,0),(7,-7),(9,2)} h(x)=3x+14 find: g^-1(0)=? h^-1(x)=? (h^-1oh)(-5)=?

OpenStudy (anonymous):

please i have no idea where to start

OpenStudy (cwrw238):

does h^-1 mean the inverse of h?

OpenStudy (anonymous):

i dont know thats all it say's

OpenStudy (cwrw238):

if so let y = 3x + 14 so x =[ y - 14] / 3 so h^-1 x = (x - 14) /3

OpenStudy (cwrw238):

i'm pretty sure thats right - im a bit rusty in functions

OpenStudy (anonymous):

thank you! i dont know but something about the way you put that answer made it make sense.

OpenStudy (cwrw238):

as for g the first numbers in brackets are the domain of g and second ones are the range domain means the values of x and range are the values of g(x)

OpenStudy (cwrw238):

(0,7) means if x = 0 then g(x) = 7. for the inverse then 7 is in domain and this maps to 0 in the range. so i think g^-1(0) = 7

OpenStudy (cwrw238):

i have no idea what (h^-1oh)(-5) means. its the oh bit that fools me

OpenStudy (anonymous):

yeah i finally figured it out it was inverse of h of h where h=-5